Please help to answer these question -

  1. The main difference between a sales tax and an excise tax is who actually bears the burden of the tax.

                              True

                              False

  1. If my price rises by $10 and my quantity falls by 10, my elasticity must be

                              1

                              -1

                              -10

                              Not enough information

  1. If my price increases from $100 to $110 and my quantity falls by 100, my elasticity is

                              1

                              -1

                              -10

                              Not enough information
Answer #1

1. The Sales Tax and Excise Duty can be distinguished in terms of the real money burden of these taxes. Due to shifting of tax a tax may be imposed on one person and it may be paid by another person. In other words the impact of a tax is on one person and the incidence on the other person. Impact is on the person on whom the tax is imposed. The incidence of a tax is on the person who actually bears the tax burden or who pay the tax.

Excise duty is the tax on the production of goods while the sales tax on the sales of the goods. In excise duty the burden is on the manufacturer who pays the tax. But in sales tax the burden of tax is on the ultimate consumer. Thus in terms of the money burden we can classify the excise duty and sales tax in to two different categories.

Answer: True

2. The elasticity of demand is the extent of change in quantity demand due to price change. In other words the elasticity measures at what rate the demand changes in response to a change in price. The elasticity can be calculated by dividing the proportionate change in demand with the proportionate change in price. If the percentage change in quantity demand and price is the same, the elasticity quotient is 1. The 10% change in price leads to 10% change in quantity demand. Then 10/10=1.

Answer: 1

If the price rise by $100 to $110 and the quantity demanded decrease by 100. Then the elasticity is change in quantity demand divided by change in price i.e 100/10=

1. It can be further explained as the initial price is $100 and current price is $110, suppose the initial quantity demanded is 200 and the current demand is 100. Then the change is price is $110-$100=10. The change in demand is 200-100=100. Then elasticity is 100/10=1.

Answer:1.